meganmoney543:

Write the equation, in general form, of the line that passes through the given points. a.y - x + 1 = 0 b.x - y + 1 = 0 c.-x - y + 1 = 0 d.x + y + 1 = 0

meganmoney543:

1 attachment
Hewoo:

What do you think the answer is?

meganmoney543:

idk but i think it might be c

Hewoo:

Ok why do you think C

meganmoney543:

idk i was jus guessing

meganmoney543:

jus being honest

meganmoney543:

i think im wrong because none of the points are negative

surjithayer:

y=x-1 y=x+1 y=-x+1 y=-x-1 find the eq. of line through points (1,2) and (4,5) slope\[=\frac{5-2}{4-1}=\frac{3}{3}=1\] eq. of line is y-2=1(x-1) y-2=x-1 y=x-1+2 y=x+1 so B

meganmoney543:

thank you so much i am very appreciative of ur help here is a medal